OpenStudy (anonymous):

Use the method of substitution to solve the system of linear equations. 3a=7-b 3a=6-b

OpenStudy (allank):

Hmm...that system does not seem possible. If 3a=7-b, then 3a cannot = 6-b. Proof: Let's make b the subject of the first equation: b=7-3a Then substitute it in the second equation: 3a=6-(7-3a) 3a=6-7+3a 0=-1

OpenStudy (allank):

Or another way to think about it, since 3a=7-b, and 3a=6-b, then that means 7-b=6-b. There is no value for b that will make this true.

OpenStudy (anonymous):

So there will be no solutions since 0=-1 is not true

OpenStudy (allank):

Yes :)